# Ultrasonic Medical Tissue Imaging Using Probabilistic Inversion: Leveraging Variational Inference for Speed Reconstruction and Uncertainty Quantification

### v2

@@ 5 Breast Data Results
Figure 5: Inversion results for the breast tissue model with a tumor. (a) The true SOS distribution in breast tissue, incorporating acoustic tissue properties and a tumor. The orange dots and green diamonds indicate the locations of the sources and receivers, respectively. (b)The SOS distribution obtained from the conventional FWI. (c) The mean result of the SVGD algorithm. The magenta arrows in (a)-(c) highlight the location of the tumor. (d) The standard deviation of the SOS distribution from the SVGD algorithm.
@@
In the previous section, the effectiveness of the SVGD algorithm is validated using both linear and ring array transducers. This section applies the method to a complex and practical medical tissue scenario. An open breast tissue database is used, incorporating acoustic tissue properties and adding a tumor, as described in [42, 43]. Figure 5(a) shows the true SOS model. The imaged object represents the SOS distribution of breast tissue, with SOS values ranging from 1.46 km/s to 1.65 km/s. The spatial discretization of the model consists of 218 x 233 grid points, with a grid spacing of 0.5 mm. The setup includes 64 sources and 64 receivers arranged in an elliptical pattern around the center of the computational domain. A Ricker wavelet with a central frequency of 0.5 MHz and a duration of 0.25 ms is used as the transmitted waveform. The remaining parameters follow the same setup as in Figure 2.
@@
In the inversion process, the initial model for FWI is set to a constant SOS of $1.5$ km/s. For the SVGD algorithm, different initial particles are generated by adding Gaussian white noise to the initial model. The prior distribution follows the Gaussian distribution (37) with $\mu_{i}=1.5$ km/s and $\sigma_{i}=1$ km/s. Figures 5(b) and (c) show the SOS distributions obtained by FWI and SVGD, respectively. Both methods successfully reconstruct the breast tissue information, identifying structures such as the skin, fat, stroma, and tumor. Figure 5(d) measures the imaging uncertainty of the SVGD mean SOS. The uncertainty varies across different tissues in the breast, with most values smaller than 0.023 and a mean uncertainty of approximately 0.012. The low uncertainty suggests that the SVGD mean SOS provides a reliable estimate of the breast tissue properties.
@@
Figure 6(a) shows the loss values plotted against the iteration number for both the conventional FWI and SVGD algorithms. The FWI algorithm reaches a lower loss value after $60$ inversion steps. Figures 6(b) and (c) display the relative error between the predicted and true SOS for FWI and SVGD, respectively. Within the gray dashed ellipse, the maximum relative error for FWI is $5.714\%$, with a mean relative error of $0.514\%$. For SVGD, the maximum relative error is $5.549\%$, and the mean relative error is $0.512\%$. Despite the slightly higher loss value for SVGD, it does not imply poorer SOS estimation. Both FWI and SVGD algorithms provide accurate and reliable estimates of the breast tissue SOS distribution.
The results in Figures 5-6 validate the effectiveness of using the mean value from SVGD for reconstructing breast tissue. To further assess the quality of the obtained standard deviation, SVGD is compared with the Stochastic Variational Inference (SVI) method introduced in [33]. The SVI-based FWI assumes a prior distribution as a standard normal distribution with mean $\bm{\mu}$ and covariance $\mathbf{\Sigma}$. This method estimates the pixel-wise standard deviation of the SOS distribution using a mean-field Gaussian approximation. The gradient of the L2-norm loss with respect to $\bm{\mu}$ (or $\mathbf{\Sigma}$) is computed using the adjoint-state method. The parameter $\bm{m}$ is derived from the updated standard normal distribution. When the loss converges or the iteration reaches a predefined threshold, the mean value is used for tissue reconstruction, and the standard deviation quantifies the uncertainty associated with the imaging results.
@@
Figure 7 (a) presents the mean SOS obtained from the SVI method, where both the breast tissue and the tumor are well reconstructed. However, the standard deviation shown in Figure 7 (b) fails to effectively quantify the uncertainty in the imaging results. In comparison to the uncertainty illustrated in Figure 5 (d), only the skin layer is clearly depicted in Figure 7 (b). The maximum standard deviation within the breast tissue is 0.14, with an average of 0.1. This standard deviation is much larger than that obtained from SVGD, suggesting that SVGD provides a tighter lower and upper bound for the prior distribution. From the variations in loss values over iterations shown in Figure 7 (c), it is evident that the SVI method achieves the converged SOS results. Figure 7 (d) shows the relative error between the predicted and true SOS for SVI. Inside the gray dashed circle, the maximum relative error is $6.121\%$, and the mean relative error is $0.558\%$. The relative error of SVI is slightly larger than that of conventional FWI and SVGD methods.
@@
The conventional FWI, SVGD, and SVI effectively reconstruct the breast tissue, with SVGD demonstrating the best performance among the three methods. To compare the reconstructed results in a local area, two lines are selected at $x=5.4$ cm and $y=5.5$ cm, and the differences in the inversion results are plotted. As shown in Figure 8(a), one vertical line at $x=5.4$ cm and one horizontal line at $y=5.5$ cm are used for comparison. Figure 8(b) shows the reconstructed SOS for the line at $x=5.4$ cm. The SOS obtained by SVGD (green line) is closer to the true value (red line), particularly in the regions where the SOS exceeds $1.55$ km/s, which suggests that SVGD is promising for approximating larger SOS areas in FWI imaging. Figure 8(c) presents the reconstructed SOS for the line at $y=5.5$ cm, the three methods align with the true SOS, and SVGD provides a better approximation in larger SOS regions.
@@
To analyze the uncertainty quantification obtained by the two VI-based methods, SVI and SVGD, Figures 9 (a) and (b) show the variations of the standard deviation along the two lines at $x=5.4$ cm and $y=5.5$ cm, respectively. Compared to the SVI, the variations in the standard deviation from the SVGD algorithm are more consistent with the true SOS variations, particularly in regions with larger SOS values. Additionally, the standard deviation values from SVGD are considerably smaller than those from SVI, indicating that SVGD provides a tighter lower and upper bound for the prior distribution.

### v3

@@ 3.3 Breast tissue imaging using SVGD and FWI algorithms
Figure 7: Inversion results for the breast tissue model with a tumor. (a) The true SOS distribution in breast tissue, incorporating acoustic tissue properties and a tumor. The orange dots and green diamonds indicate the locations of the sources and receivers, respectively. (b) The SOS distribution obtained from the conventional FWI. (c) The mean result of the SVGD algorithm using 20 particles. The magenta arrows in (a)-(c) highlight the location of the tumor. (d) The standard deviation of the SOS distribution from the SVGD algorithm.
@@
In the previous paragraphs, the effectiveness of the SVGD algorithm is validated using both linear and ring array transducers. To further assess the performance under complex anatomical structures, we apply SVGD to a simulated breast tissue model derived from an open database. The model incorporates realistic acoustic tissue properties and includes an embedded tumor, as described in [45, 46]. Figure 7(a) shows the true SOS model. The imaged object represents the SOS distribution of breast tissue, with SOS values ranging from 1.46 km/s to 1.65 km/s. The computational domain is discretized into $356\times 385$ grid points with a uniform spatial spacing of 0.3 mm. A total of 64 sources and 64 receivers are uniformly arranged in an elliptical pattern centered within the domain. The transmitted waveform is a Ricker wavelet with a central frequency of 0.5 MHz and a duration of 0.09 ms. The time step $\Delta t$ is determined based on the CFL condition, using the stability constant $C=0.3$, resulting in $\Delta t=5\times 10^{-5}$ ms and a total of 1801 temporal samples. The spatial discretization corresponds to approximately 10 grid points per wavelength.
@@
In the inversion process, the initial model for conventional FWI is set to a constant SOS of $1.5$ km/s, resulting in an initial squared slowness of $\bm{m}=1/\bm{c}^{2}\approx 0.444~\text{s}^{2}/\text{km}^{2}$. The true SOS varies within the range $c_{\min}=1.46$ km/s to $c_{\max}=1.65$ km/s, and the squared slowness is constrained within $m_{\min}=1/c_{\max}^{2}$ and $m_{\max}=1/c_{\min}^{2}$ during the inversion. The optimization is performed using the quasi-Newton method L-BFGS-B, implemented via the scipy.optimize.minimize function in Python, with a maximum of 50 iterations. The algorithm utilizes an internal line search procedure in which the step length is adaptively determined at each iteration. The stopping criterion is satisfied when the relative change in the objective function falls below $10^{-5}$ or when the maximum number of iterations is reached.
The SVGD initialization involves generating 20 particles by injecting Gaussian noise (mean 0, standard deviation 0.01) into the initial FWI model. The prior distribution follows the truncated Gaussian distribution (23) with $\mu_{i}=0.444~\text{s}^{2}/\text{km}^{2}$ and $\sigma_{i}=0.1~\text{s}^{2}/\text{km}^{2}$. The optimization is performed using the quasi-Newton method L-BFGS-B, with a maximum of 50 iterations. The algorithm employs an internal line search procedure, and the step length is adaptively determined. The stopping criterion is met when the objective change is below $10^{-5}$ or the iteration limit is reached.
@@
Figures 7(b) and (c) show the SOS distributions obtained by FWI and SVGD, respectively. Both methods successfully reconstruct the breast tissue information, identifying structures such as the skin, fat, stroma, and tumor. Figure 7(d) measures the imaging uncertainty of the SVGD mean SOS. The uncertainty varies across different tissues in the breast, with most values smaller than 0.01 and a mean uncertainty of approximately 0.003. The low uncertainty suggests that the SVGD mean SOS provides a reliable estimate of the breast tissue properties.
@@
Figure 8(a) shows the loss values plotted against the iteration number for both the conventional FWI and SVGD algorithms. The FWI algorithm reaches a lower loss value after 50 inversion steps, which is due to the different loss functions used by FWI and SVGD. FWI minimizes the $l_{2}$ data misfit, while SVGD minimizes the KL divergence between the unnormalized posterior and the inferred distribution. Figures 8(b) and (c) display the relative error between the predicted and true SOS for FWI and SVGD, respectively. Within the gray dashed ellipse, the maximum relative error for FWI is $1.25\%$, with a mean relative error of $0.08\%$. For SVGD, the maximum relative error is $1.10\%$, and the mean relative error is $0.09\%$. Both FWI and SVGD algorithms provide accurate and reliable estimates of the breast tissue SOS distribution.
Figure 9: Sensitivity of the SVGD algorithm to particle count in breast tissue data. (a) Mean SOS for $n=3,5,10,15$ (from left to right). (b) Standard deviation of the SOS distribution for each particle count.
Figure 9 demonstrates the sensitivity of the SVGD algorithm to the number of particles used in the inversion process. Panels (a) and (b) show the mean and standard deviation of the SOS, respectively, for particle counts of $n=3,5,10,15$ (from left to right). Even with a small number of particles, the SVGD algorithm produces reliable mean SOS results. As the number of particles increases, the standard deviation distribution becomes more refined, with clearer boundaries between different tissue types. The results suggest the advantages of the SVGD algorithm, demonstrating that the method is scalable by adjusting the particle count. Even with fewer particles, the algorithm still achieves satisfactory inversion results, highlighting its efficiency and flexibility in handling different particle sizes.
The results in Figures 7-8 validate the effectiveness of using the mean value from SVGD for reconstructing breast tissue. To further assess the quality of the obtained standard deviation, we conduct a comparative analysis of SVGD against two Bayesian inference methods, namely SVI [34] and Metropolis-Hastings MCMC [47, 48]. This comparison enables a more rigorous evaluation of SVGD’s capability for uncertainty quantification.
3.4 Uncertainty Quantification via SVI-Based FWI and Metropolis-Hastings MCMC
@@
The SVI-based FWI is a probabilistic approach that leverages stochastic variational inference to approximate the posterior distribution of SOS. SVI assumes a mean-field Gaussian prior that is spatially uncorrelated, with an initial mean $\tilde{\bm{\mu}}$ and standard deviation $\tilde{\bm{\sigma}}$. To enhance the accuracy and stability of gradient estimation, the method draws $n$ samples $\bm{c}_{i}^{t}$ at each iteration $t$ from the current Gaussian prior $\mathcal{N}(\tilde{\bm{\mu}}^{t},\tilde{\bm{\sigma}}^{t})$. The mean and standard deviation of the prior are updated by [34]
$\tilde{\bm{\mu}}^{t+1}=\tilde{\bm{\mu}}^{t}+\frac{1}{n}\sum_{i=1}^{n}\mathbf{\Delta}_{i}^{t},~~\tilde{\bm{\sigma}}^{t+1}=\tilde{\bm{\sigma}}^{t}+\frac{1}{n}\sum_{i=1}^{n}\bm{\eta}_{i}^{t}\cdot\mathbf{\Delta}_{i}^{t},$ (25)
where $\bm{\eta}_{i}^{t}\sim\mathcal{N}(0,\mathbf{I})$ introduces stochastic perturbations. The term $\mathbf{\Delta}_{i}^{t}$ denotes the adjoint-state gradient of the L2-norm data misfit with respect to sample $\bm{c}_{i}^{t}$. Upon convergence or after a predefined number of iterations, the mean value is used to reconstruct the tissue properties, and the standard deviation quantifies voxel-wise uncertainty.
For the SVI-based inversion, the initial prior distribution is modeled as a truncated Gaussian with mean $\tilde{\mu}_{i}^{0}=1.5~\text{km}/\text{s}$ and standard deviation $\tilde{\sigma}_{i}^{0}=0.1~\text{km}/\text{s}$. The prior distribution is iteratively refined using the SVI update rule Eq. (25). At each iteration, 20 samples $\bm{c}_{i}^{t}$ are generated from the current prior distribution. The optimization follows a steepest descent scheme with a maximum of 100 iterations. The step size is initially set to $5\times 10^{-3}$ and is reduced by a factor of 0.8 every 20 iterations. The inversion terminates when either the relative change in the objective function falls below $10^{-5}$ or the iteration limit is reached.
@@
Figure 10(a) presents the mean SOS obtained from the SVI method, where both the breast tissue and the tumor are well reconstructed. However, the standard deviation shown in Figure 10(b) fails to effectively quantify the uncertainty in the imaging results. In comparison to the uncertainty illustrated in Figure 7(d), only the skin layer is clearly depicted in Figure 10(b). The maximum standard deviation within the ROI is 0.14, with an average of 0.1, which is larger than that obtained from SVGD. From the variations in loss values over iterations shown in Figure 10(c), it is evident that the SVI method achieves the converged SOS results. Figure 10(d) shows the relative error between the predicted and true SOS for SVI. Inside the gray dashed circle, the maximum relative error is $5.054\%$, and the mean relative error is $0.396\%$. The relative error of SVI is larger than that of conventional FWI and SVGD methods.
The standard deviation of the posterior distribution reflects the intrinsic uncertainty of inverse problems. It primarily depends on the choice of the prior distribution and the level of observational noise. Therefore, different algorithms should not produce significantly different estimates of posterior uncertainty. To provide a reliable reference for uncertainty quantification, we employ the Metropolis-Hastings MCMC (MH-MCMC) method to sample from the posterior distribution of the model parameters.
The MH-MCMC algorithm constructs a Markov chain whose stationary distribution converges to the target posterior distribution after sufficient iterations. The algorithm starts by drawing a candidate sample $\bm{m}^{\prime}$ from a proposal distribution $q(\bm{m}^{\prime}\mid\bm{m}^{t})$, which denotes the probability of $\bm{m}^{\prime}$ given the current state $\bm{m}^{t}$. The candidate is accepted with probability $\alpha$, defined as
$\alpha=\min\left\{1,\frac{p(\bm{m}^{\prime}\mid\bm{d}^{\text{obs}})\,q(\bm{m}^{t}\mid\bm{m}^{\prime})}{p(\bm{m}^{t}\mid\bm{d}^{\text{obs}})\,q(\bm{m}^{\prime}\mid\bm{m}^{t})}\right\},$ (26)
where $p(\bm{m}^{t}\mid\bm{d}^{\text{obs}})$ denotes the posterior probability, and $q(\bm{m}^{t}\mid\bm{m}^{\prime})$ is the reverse proposal distribution. If accepted, the Markov chain moves to the candidate state, i.e., $\bm{m}^{t+1}=\bm{m}^{\prime}$. Otherwise, it remains at the current state, $\bm{m}^{t+1}=\bm{m}^{t}$. This process is repeated until a specified number of iterations is reached or convergence is achieved.
To enhance sampling efficiency and acceptance rate, the Metropolis-adjusted Langevin Algorithm (MALA) is incorporated into the MH-MCMC framework. MALA leverages gradient-informed proposals to improve sampling directionality and mitigate the inefficiency of random-walk behavior in traditional MH-MCMC algorithms. The proposal distribution for generating candidate samples $\bm{m}^{\prime}$ is [47, 48]
$\bm{m}^{\prime}=\bm{m}^{t}+\frac{\epsilon_{t}^{2}}{2}\nabla\log p(\bm{m}^{t}\mid\bm{d}^{\text{obs}})+\epsilon_{t}\bm{\eta}_{t},$ (27)
where $\bm{\eta}_{t}\sim\mathcal{N}(0,\mathbf{I})$ is a $d$-dimensional standard Gaussian random variable. Each update in Eq. (27) combines a gradient-driven drift term, which pulls particles toward high-probability regions of the posterior, with a stochastic diffusion term that preserves exploration and ensures ergodicity. The step size $\epsilon_{t}$ determines how much influence the gradient term and the noise term have on the candidate sample. This combination enables more efficient exploration of the posterior distribution in high-dimensional inverse problems.
In practical implementations of the MH-MCMC method, an initial set of samples, referred to as the burn-in period, is typically discarded to allow the Markov chain to converge to its stationary distribution. To accelerate convergence and ensure high-quality initialization in posterior sampling of breast tissue data, a two-stage strategy is employed. During the burn-in stage, 20 chains are initialized from a normal distribution with a mean of $1/1.5^{2}$ (assuming the SOS in pure water is 1.5 km/s) and a standard deviation of 0.01. To emphasize the accurate estimation of posterior uncertainty, the prior distribution $p(\bm{m})$ is also the truncated Gaussian, centered at the reciprocal of the squared true SOS, with a standard deviation of 0.1. The MALA proposal Eq. (27) is used to update each of the 20 chains. The initial step size is set to $\epsilon_{t}=1\times 10^{-2}$, with a gradual reduction in step size over 100 iterations, and all samples from this stage are discarded to eliminate initialization bias [47]. In the subsequent MH-MCMC sampling stage, 100 iterations are performed using a smaller step size of $\epsilon_{t}=2\times 10^{-4}$. At each iteration, a candidate sample is generated and evaluated using the Metropolis-Hastings acceptance probability $\alpha$. If the candidate is accepted, it replaces the current sample; otherwise, the chain remains at the current state.
Figure 11: MH-MCMC sampling results for breast tissue SOS reconstruction. (a) Acceptance rates of 20 chains. (b) Posterior mean of the SOS obtained from 2000 samples. (c) Posterior standard deviation of the SOS. (d) Relative error of the reconstructed SOS.
@@
After performing 100 iterations of MH-MCMC sampling, the algorithm generates 100 samples for each of the 20 chains, resulting in a total of 2000 realizations of the SOS. Figure 11(a) shows the acceptance rates for the 20 chains, ranging from 0.52 to 0.83, with an average of 0.7295, indicating that most candidate samples are accepted. The posterior mean and standard deviation, calculated from the 2000 samples, are presented in Figures 11(b) and (c), respectively. As shown in Figure 11(c), the uncertainty map within the ROI highlights the delineation of tissue boundaries. Figure 11(d) presents the relative error of the reconstructed SOS, which reaches a maximum of 1.850% and averages 0.248% within the ROI.
The conventional FWI, SVGD, MH-MCMC, and SVI effectively reconstruct breast tissue structures. To evaluate local reconstruction accuracy, two cross-sectional lines at $x=5.4$ cm and $y=5.5$ cm are selected for comparison, as illustrated in Figure 12(a). Figure 12(b) shows the reconstructed SOS for the line at $x=5.4$ cm. The results obtained by conventional FWI, SVGD, and MH-MCMC closely match the true value, whereas SVI yields slightly less accurate estimates in regions of higher SOS. Figure 12(c) presents the results along the horizontal line at $y=5.5$ cm, where SVGD again demonstrates comparable accuracy to conventional FWI. In addition to its comparable accuracy, SVGD provides uncertainty quantification and is computationally more efficient than MH-MCMC, making it a promising method for SOS evaluation in FWI imaging.
@@
To evaluate the quality of uncertainty quantification produced by the two variational inference methods, SVI and SVGD, Figures 13(a) and (b) present the standard deviation profiles along the lines at $x=5.4$ cm and $y=5.5$ cm, respectively. Compared to SVI, SVGD yields standard deviation estimates that better capture the variations of the true SOS, particularly in regions with higher SOS values, and exhibits strong agreement with the MH-MCMC reference. This observation is further supported by quantitative analysis using the Pearson correlation coefficient. Along the line at $x=5.4$ cm, the correlation between SVGD and MH-MCMC reaches 0.65, in contrast to only 0.01 for SVI. Besides, along the line at $y=5.5$ cm, the correlation between SVGD and MH-MCMC is 0.69, whereas the correlation between SVI and MH-MCMC is 0.07. The results demonstrate that SVGD produces uncertainty estimates more consistent with the MH-MCMC benchmark than SVI, underscoring its capability for reliable uncertainty quantification.
3.5 Computational and memory cost analysis
When applying the SVGD update rule Eq. (15) to inversion imaging, both computational and memory costs should be considered. The computational cost is primarily determined by the number of wave equation solves required per iteration. In each iteration, updating all particles $\{\bm{m}_{j}^{t}\}_{j=1}^{n}$ requires computing the gradient of the log-posterior $\nabla_{\bm{m}_{i}^{t}}\log p\left(\bm{m}_{i}^{t}\mid\bm{d}^{\mathrm{obs}}\right)$, which consists of the likelihood gradient $\nabla_{\bm{m}}\log p\left(\bm{d}^{\mathrm{obs}}\mid\bm{m}\right)$ and the prior gradient $\nabla_{\bm{m}}\log p\left(\bm{m}\right)$. The likelihood gradient is computed using the adjoint-state method, requiring one forward and one adjoint wave equation solve for each emitting transducer. In contrast, the prior gradient, derived analytically from the truncated Gaussian prior, incurs negligible computational cost. Consequently, each particle requires $2N_{s}$ wave equation solves per iteration, where $N_{s}$ denotes the number of emitting transducers. For $n$ particles, the total cost per iteration is $2nN_{s}$ wave equation solves. In comparison, standard deterministic FWI requires only $2N_{s}$ wave equation solves per iteration. However, SVGD enables posterior distribution approximation and uncertainty quantification, with computational cost scaling linearly with $n$. In our implementation, we typically use $n=20$ particles, resulting in $40N_{s}$ wave equation solves per iteration. Empirical results, such as those shown in Figures 5 and 10, demonstrate that using fewer particles (e.g., $n=10$ or $15$) still yields high-quality inversion results. If moderate uncertainty resolution is sufficient, even 5 particles can provide acceptable reconstructions, thus allowing a trade-off between computational efficiency and the fidelity of uncertainty quantification.
The computational cost of the SVI-based FWI method is primarily determined by the number of wave equation solves required per iteration. At each iteration, $n$ samples $\{\bm{c}_{i}\}_{i=1}^{n}$ are drawn from the current Gaussian distribution. For each sample, the algorithm performs one forward and one adjoint wave equation solve for each emitting transducer to evaluate the misfit and compute its gradient via the adjoint-state method. Therefore, the total number of wave equation solves per iteration is $2nN_{s}$. Although both SVGD and SVI incur the same wave solver cost per iteration, SVGD explicitly evolves a set of interacting particles and leverages kernelized gradient updates, whereas SVI maintains and updates only the mean and covariance of the approximate posterior.
Table 1: Comparison of computational cost over 100 iterations ($N_{s}=64$).
Method Particles ($n$)
Wave Solves
per Iteration
Total Wave
Solves (100 iters)
Time per
Iteration (s)
Total Time
FWI $2N_{s}=128$ 12,80012{,}800 26.04 43.4
SVGD 20 $2nN_{s}=2560$ 256,000256{,}000 528.6 881
SVI 20 $2nN_{s}=2560$ 256,000256{,}000 530.9 884.8
To compare the computational efficiency of different inversion methods, we apply conventional FWI, SVGD, and SVI to the breast tissue model shown in Figure 7(a). For a fair comparison, all methods adopt the steepest descent optimization algorithm, with identical configurations for both forward and inverse modeling. Experiments are conducted on a high-performance computing (HPC) cluster running Ubuntu 18.04.6 LTS, equipped with four Intel Xeon E7-4830 v4 processors (112 threads in total), 503 GB of RAM, and Linux kernel version 5.4.0. The Devito software package (https://github.com/devitocodes/devito) is used to compute the forward wavefield and the gradient via the adjoint-state method. Computational times are averaged over 100 inversion iterations. As summarized in Table 1, the per-iteration computational time of SVGD and SVI is approximately $n$ times that of FWI, with runtimes of 528.6 s and 530.9 s, respectively, compared to 26.04 s for FWI. Note that the reported computational times already include the effect of $N_{s}=64$ emitting transducers.
To estimate the memory requirements of the proposed SVGD-based FWI method, we analyze the dominant components per iteration for $\bm{m}\in\mathbb{R}^{d}$. First, particle storage requires maintaining $n$ particles, leading to a memory cost of $\mathcal{O}(nd)$. Second, the gradient of the log-posterior for each particle must be stored, also incurring a cost of $\mathcal{O}(nd)$. Third, computation of the kernel matrix $\mathbf{K}\in\mathbb{R}^{n\times n}$ and its gradient $\nabla_{\bm{m}_{i}}k(\bm{m}_{i},\bm{m}_{j})\in\mathbb{R}^{d}$ for all pairs $(i,j)$ results in a memory cost of $\mathcal{O}(n^{2}d)$. Fourth, each particle must store forward and adjoint wavefields across a space-time grid. The dominant term for each wavefield is $\mathcal{O}(dT)$, yielding a total memory cost of $\mathcal{O}(nN_{r}dT)$ across all particles, where $N_{r}$ is the number of receiving transducers. Since the method employs single-source emission, the number of emitting transducers $N_{s}$ does not significantly affect memory storage for the wavefields. Other temporary variables, such as misfit values or prior terms, contribute negligibly to the overall cost. Summing the dominant terms, the total memory requirement is $\mathcal{O}(nd(1+n+T+N_{r}))$, which scales linearly with $d$, $T$, and $N_{r}$, and quadratically with $n$. For moderate values of $n$ (e.g., 10–20), the method remains memory-efficient relative to conventional ensemble-based Bayesian sampling approaches.